# Analyzing the Effect of Linguistic Similarity on Cross-Lingual Transfer: Tasks and Experimental Setups Matter

### v1

@@ 4.4 Practical implications: Choosing transfer languages based on simple heuristics
Src = best src lang for the target in another experiment POS deV POS LAS UAS to. b. to. tr. to. m. POS deV 304 406 506 610 712 606 POS 609 205 305 911 1012 1614 LAS 709 102 102 1114 1013 1814 UAS 810 305 102 1215 1012 1714 topics b. 1211 1011 1112 1313 307 1915 topics tr. 1009 808 808 909 306 1813 topics m. 506 403 505 505 707 607
@@
Table 1: Mean performance loss if picking a source language solely based on a given metric (top) or based on its transfer performance in another experiment (bottom). Performance loss in percentage points. Subscripts = standard deviations. “Word*” is word overlap for the UD tasks, trigram overlap for the n-gram topic classification models, and subword token overlap for the mBERT-based MLP.
So far we have investigated correlations on a global level. In this section we investigate how these findings can be used in practice: Do the overall correlation patterns also apply when picking a transfer language for a given target language according to simple heuristics derived from the global patterns?
Based on similarity measures
@@
Take-away 1: The general correlation patterns still apply when picking a transfer language for a given target language. The patterns for the most strongly correlated similarity measures are similar: e.g., parsing performance is most strongly correlated with syntactic similarity, and picking the syntactically most similar language as the source also results in the lowest performance loss for parsing (Table 1, top).
@@
However, some of the less strong predictors in the global correlations are nearly as good for picking a source language, e.g., genetic and lexical similarity for parsing. We hypothesize that this is due some linguistic measures being more correlated when similarity is high. E.g., very closely related languages tend to be syntactically similar, even if the overall correlation between genetic and syntactic similarity is only moderately high due the existence of languages that are unrelated but share syntactic similarities or that are (more distantly) related and syntactically dissimilar (Figure 6 in §G).
Conversely, character overlap is a worse measure for selecting a source language than the overall correlations would suggest. This is likely due to transfer between languages with low character overlap generally working poorly, while high overlap does not guarantee good transfer (e.g., Figure 7 in Appendix §G).
@@
We additionally simulate a set-up where a researcher has data in multiple source language candidates at their disposal and can afford training and comparing a small selection of models. We select the top three most similar languages according to each measure and take the highest transfer score produced by any of them (Table 10 in §G). The same trends still hold as when picking only one source language candidate. However, the overall results are much better, and the gaps between the performance losses between the different measures become smaller. Take-away 2: Comparing the results of multiple top source language candidates often yields better results than only considering the most similar source language.
Based on the results of another experiment
@@
Take-away: Choosing a source language based on the results for a similar task with similar input representations leads to only small losses. These losses are often even smaller than when selecting source languages based on well-chosen similarity measures. However, using other input representations weakens the predictive effect somewhat: e.g., for picking source languages for POS tagging with UDPipe, it is worse to choose based on de Vries et al.’s XLM-R results than to choose based on UDPipe’s parsing results. Picking transfer languages based on other tasks leads to bigger losses. Topics-mbert is an especially poor predictor – this method likely suggests a fairly arbitrary mBERT language, which might or might not be similar to the target language.
@@ 5 Conclusion
We investigated how linguistic and dataset level features impact cross-lingual transfer, and find that the most important features differ across tasks and input representations. If we zoom in on a language level, we find that a feature’s predictive performance on one task transfers relatively well to another similar task. Concretely, for practitioners, this means that one should ideally pick a transfer language based on strongly predictive similarity measures for a task. However, if these are not available, selecting transfer languages based on the similarity measures that are predictive for a similar task is often a safe choice. Future work should investigate to what extend these patterns hold across even more languages, tasks, and NLP models.
@@
Within linguistics, there is a rich literature on defining similarity measures Borin (2013). We chose to include similarity measures that are commonly used in NLP studies, and we adapt them where necessary. However, we acknowledge that there are many other similarity measures that are interesting to investigate.
@@ Appendix A Languages and Datasets
Tables 2 and 3 show the languages and datasets included in our experiments. Their geographic distribution is pictured in Figure 1.
@@ A.2 Excluded SIB-200 languages
We exclude three Arabic dialects whose sentences in SIB-200 are almost identical to the Modern Standard Arabic (arb_Arab) sentences: ars_Arab, acm_Arab, acq_Arab. This is a known issue for FLORES-200, from which SIB-200 is derived. 999github.com/openlanguagedata/flores/issues/8
@@ Appendix B Resources Used
Resource Resource URL Universal Dependencies 2.14 Zeman et al. (2024) hdl.handle.net/11234/1-5502 SIB-200 Adelani et al. (2024) huggingface.co/datasets/Davlan/sib200 Grambank v1.0.3 Skirgård et al. (2023a) zenodo.org/records/7844558 Glottolog v5.0 Hammarström et al. (2024) zenodo.org/records/10804582 ASJP v17 Wichmann et al. (2016) asjp.clld.org Jäger (2018a) osf.io/cufv7 lang2vec v1.1.6 (Littell et al., 2019) github.com/antonisa/lang2vec UDPipe 2 (2.12) (Straka, 2023) https://ufal.mff.cuni.cz/udpipe/2/models mBERT base uncased Devlin et al. (2019) huggingface.co/google-bert/bert-base-multilingual-uncased lme4 Bates et al. (2015) github.com/lme4/lme4 scikit-learn 1.5 Pedregosa et al. (2011) scikit-learn.org uroman 1.3.1.1 Hermjakob et al. (2018) github.com/isi-nlp/uroman GlotScript Kargaran et al. (2024) github.com/cisnlp/GlotScript SciPy Virtanen et al. (2020) scipy.org
@@
Lang2vec Littell et al. (2017) contains information on syntax, phonology, and phoneme inventories from WALS Dryer and Haspelmath (2013), SSWL Collins and Kayne (2011), Ethnologue Lewis et al. (2015), and PHOIBLE Moran et al. (2014). SIB-200 Adelani et al. (2024) is an annotated subset of FLORES-200 (NLLB Team et al., 2022), which in turn builds on previous versions and extensions of FLORES (Goyal et al., 2022; Guzmán et al., 2019; Doumbouya et al., 2023; AI4Bharat et al., 2023). UDPipe 2 (Straka, 2023) was trained on Universal Dependencies 2.12.101010http://hdl.handle.net/11234/1-5150 Its input representations include the last four layers of base-size uncased mBERT Devlin et al. (2019). This project uses the universal romanizer software ‘uroman’ written by Ulf Hermjakob, USC Information Sciences Institute (2015-2020).
@@ E.2 Robustness across datasets of the same language
Most treebank pairs of the same language also show very similar transfer patterns. Again, most correlations are close to 1.121212For the training treebanks, we see correlations ($r$) below $0.8$ only for the following pairs: the two German training sets (hdt & gsd: $r=0.75$), the two Persian ones (perft & seraji: $r=0.73$), and the two Korean ones (gsd & kaist: $r=0.72$). For the test datasets this applies to some of the six Latin treebanks (circse & udante: $r=0.79$, circse & ittb: $r=0.79$), several of the eight Turkish datasets (atis & tourism: $r=0.55$, pud & tourism: $r=0.73$, penn & tourism: $r=0.74$, boun & tourism: $r=0.74$, gb & tourism: $r=0.78$, kenet & tourism: $r=0.79$), several of the seven Mandarin Chinese treebanks (hk & patentchar: $r=0.77$, beginner & patentchar: $r=0.78$, patentchar & pud: $r=0.79$, cfl & patentchar: $r=0.79$, gsd & patentchar: $r=0.79$), two of the four Old East Slavic treebanks (birchbark & rnc: $r=0.75$), the two Akkadian treebanks (pisandub & riao: $r=0.75$), and both Sanskrit treebanks (ufal & vedic: $r=0.57$).
@@
E.3 Transfer across scripts in SIB-200
Eight of the languages in SIB-200 come in two script versions each. The writing systems are completely distinct, except for Mandarin Chinese, which has versions in traditional and simplified characters. For topics-base, transfer between the two scripts of a language is unsurprisingly always much lower than the performance on the same script (with accuracies between 8.3% and 25.9% for in-language cross-script transfer – excluding the Mandarin Chinese entries, for which cross-script accuracy is up to 64.7% – vs. within-language, within-script accuracies between 60.8% and 75.9% for the same languages). Although transfer between the transliterated versions of the datasets works better (between 14.2% and 50.5%), the accuracies are still much lower than the within-language, within-original-script accuracies (between 58.3% and 74.0%). This is likely due to different transliteration conventions for different writing systems (and due to missing vowels in the transliterated versions of abjads). The patterns are also similar for topics-mbert, with within-language, cross-script scores between 8.3% and 45.6%, compared to within-language, within-script accuracies between 42.2% and 79.9%.
E.4 Effects of transliteration
@@ F.1 Correlations between task results
Table 8 shows the correlations between task results. It is described in §4.1.2. The correlations across different task types only involve the 55 training and 84 test languages that appear both in UD and in SIB-200 (54 and 82 in the comparisons with the transliterated SIB-200 data).
@@ F.3 Correlations between NLP results and similarity measures
Tables 12 and 13 show the correlations between POS/parsing scores and the similarity measures for each test language. Tables 14, 15, and 16 show the same for the topic classification experiments. Because these tables take up a lot of space, they are placed at the very end of the appendix.
Appendix G Implications for Choosing Transfer Languages
@@
This section provides additional context for §4.4. Table 10 shows the mean performance loss when choosing source languages based on their similarity to the target language (or based on their training data size). Table 10 shows the loss when choosing the source languages that were the best sources in another cross-lingual transfer task.

### v2

@@ 5.2.1 Should I pick the most similar language according to one similarity measure?
Src = best src lang for the target in another experiment POS deV POS LAS UAS to.-b. to.-t. to.-m. POS deV 304 406 506 610 712 606 POS 609 205 305 911 1012 1614 LAS 709 102 102 1114 1013 1814 UAS 810 305 102 1215 1012 1714 topics b. 1211 1011 1112 1313 307 1915 topics tr. 1009 808 808 909 306 1813 topics m. 506 403 505 505 707 607
@@
Table 3: Mean performance loss if picking a source language solely based on a given metric (top) or based on its transfer performance in another experiment (bottom). Performance loss in percentage points. Subscripts = standard deviations; “word*” = wor, tri, swt; deV = de Vries et al. (2022).
@@
The patterns for the most strongly correlated similarity measures are similar – choosing a source language based on the pertinent similarity measure incurs relatively small losses. E.g., parsing performance is most strongly correlated with syntactic similarity, and picking the syntactically most similar language as the source also results in the lowest performance loss for parsing (Table 3, top). Our findings suggest choosing a syntactically similar language for POS tagging and dependency parsing, and a dataset with high trigram overlap for the $n$-gram-based topic classification experiments.
@@
However, some of the less strong predictors in the global correlations are nearly as good for picking a source language, e.g., genetic and lexical similarity for parsing. We hypothesize that this is due some linguistic measures being more correlated when similarity is high. [^14] E.g., very closely related languages tend to be syntactically similar, even if the overall correlation between genetic and syntactic similarity is only moderately high due the existence of languages that are unrelated but share syntactic similarities or that are (more distantly) related and syntactically dissimilar (Figure 6 in §G). Conversely, character overlap is a worse measure for selecting a source language than the overall correlations would suggest. This is likely due to transfer between languages with low character overlap generally working poorly, while high overlap does not guarantee good transfer (e.g., Figure 7 in Appendix §G).
@@
We additionally simulate a set-up where a researcher has data in multiple source language candidates at their disposal and can afford training and comparing a small selection of models. We select the top three most similar languages according to each measure and take the highest transfer score produced by any of them (Table 12 in §G). The same trends still hold as when picking only one candidate. However, the overall results are much better, and the gaps between the performance losses between the different measures become smaller. Comparing the results of multiple top source language candidates often yields better results than only considering the most similar one.
5.2.2 Should I choose a source language based on another experiment?
@@
For the tasks included in our experiments, choosing a source language based on the results for a similar task with similar input representations leads to only small losses: POS tagging results serve as a good predictor for parsing and vice versa, but not for topic classification. Similarly, the results for some topic classification settings are good predictors for each other, but not for POS tagging and parsing. These losses are often even smaller than when selecting source languages based on similarity measures (§5.2.1, Table 3, top).
However, using other input representations weakens the predictive effect somewhat: e.g., for picking source languages for POS tagging with UDPipe, it is worse to choose based on de Vries et al.’s XLM-R results than to choose based on UDPipe’s parsing results. Topics-mbert is an especially poor predictor – this method likely suggests a fairly arbitrary mBERT language, which might or might not be similar to the target.
@@ 6 Conclusion
We investigated how linguistic and dataset level similarities impact cross-lingual transfer, and find that the most predictive similarity measures differ across tasks and input representations. For practitioners working on the tasks included in our study, we recommend choosing a source language based on a pertinent similarity measure, or, if results for an extensive transfer experiment involving a similar task and similar input representations are available, based on the transfer results in that experiment. Future work should investigate to what extent these patterns hold across even more languages, tasks, and NLP models.
@@
We do not consider model- or tokenization-specific biases that may have different effects based on the word order White and Cotterell (2021), morphology Park et al. (2021), or writing system and orthography Sproat and Gutkin (2021).
@@
Within linguistics, there is a rich literature on defining similarity measures (cf. Borin, 2013). We chose to include similarity measures that are commonly used in NLP studies, and we adapt them where necessary. However, we acknowledge that there are many other similarity measures that are interesting to investigate.
@@ Appendix A Languages and Datasets
Tables 4 and 5 show the languages and datasets included in our experiments. Their geographic distribution is pictured in Figure 1.
@@ A.2 Excluded SIB-200 languages
We exclude three Arabic dialects whose sentences in SIB-200 are almost identical to the Modern Standard Arabic (arb_Arab) sentences: ars_Arab, acm_Arab, acq_Arab. This is a known issue for FLORES-200, from which SIB-200 is derived.151515github.com/openlanguagedata/flores/issues/8
@@ Appendix B Resources Used
Resource Resource URL Universal Dependencies 2.14 Zeman et al. (2024) hdl.handle.net/11234/1-5502 SIB-200 Adelani et al. (2024) huggingface.co/datasets/Davlan/sib200 Grambank v1.0.3 Skirgård et al. (2023) zenodo.org/records/7844558 Glottolog v5.0 Hammarström et al. (2024) zenodo.org/records/10804582 ASJP v17 Wichmann et al. (2016) asjp.clld.org Jäger (2018a) osf.io/cufv7 lang2vec v1.1.6 (Littell et al., 2019) github.com/antonisa/lang2vec UDPipe 2 (2.12) (Straka, 2023) https://ufal.mff.cuni.cz/udpipe/2/models mBERT base uncased Devlin et al. (2019) huggingface.co/google-bert/bert-base-multilingual-uncased lme4 Bates et al. (2015) github.com/lme4/lme4 scikit-learn 1.5 Pedregosa et al. (2011) scikit-learn.org uroman 1.3.1.1 Hermjakob et al. (2018) github.com/isi-nlp/uroman GlotScript Kargaran et al. (2024) github.com/cisnlp/GlotScript SciPy Virtanen et al. (2020) scipy.org
@@
Lang2vec Littell et al. (2017) contains information on syntax, phonology, and phoneme inventories from WALS Dryer and Haspelmath (2013), SSWL Collins and Kayne (2011), Ethnologue Lewis et al. (2015), and PHOIBLE Moran et al. (2014). SIB-200 Adelani et al. (2024) is an annotated subset of FLORES-200 (NLLB Team et al., 2022), which in turn builds on previous versions and extensions of FLORES (Goyal et al., 2022; Guzmán et al., 2019; Doumbouya et al., 2023; AI4Bharat et al., 2023). UDPipe 2 (Straka, 2023) was trained on Universal Dependencies 2.12.161616http://hdl.handle.net/11234/1-5150 Its input representations include the last four layers of base-size uncased mBERT Devlin et al. (2019). This project uses the universal romanizer software ‘uroman’ written by Ulf Hermjakob, USC Information Sciences Institute (2015-2020).
@@ E.2 Robustness across datasets of the same language
Most treebank pairs of the same language also show very similar transfer patterns. Again, most correlations are close to 1.181818For the training treebanks, we see correlations ($r$) below $0.8$ only for the following pairs: the two German training sets (hdt & gsd: $r=0.75$), the two Persian ones (perft & seraji: $r=0.73$), and the two Korean ones (gsd & kaist: $r=0.72$). For the test datasets this applies to some of the six Latin treebanks (circse & udante: $r=0.79$, circse & ittb: $r=0.79$), several of the eight Turkish datasets (atis & tourism: $r=0.55$, pud & tourism: $r=0.73$, penn & tourism: $r=0.74$, boun & tourism: $r=0.74$, gb & tourism: $r=0.78$, kenet & tourism: $r=0.79$), several of the seven Mandarin Chinese treebanks (hk & patentchar: $r=0.77$, beginner & patentchar: $r=0.78$, patentchar & pud: $r=0.79$, cfl & patentchar: $r=0.79$, gsd & patentchar: $r=0.79$), two of the four Old East Slavic treebanks (birchbark & rnc: $r=0.75$), the two Akkadian treebanks (pisandub & riao: $r=0.75$), and both Sanskrit treebanks (ufal & vedic: $r=0.57$).
@@
E.3 Effect of writing system
Transfer between datasets with the same writing systems generally works better than between different scripts, however this is in part due to the language selection rather than the scripts themselves. Two-sample Kolmogorov–Smirnov tests indicate that the results within vs. across scripts come from different distributions (p-values all <0.0001; statistics: 0.51 for topics-base, 0.26 for topics-mbert, 0.16 for POS accuracy, 0.18 for LAS). However, the results of topics-translit for datasets that were in the same vs. different scripts before transliteration also come from different distributions (p <0.0001, statistic: 0.37), indicating that at least for SIB-200, the combinations of languages usually associated with the scripts alone already make an important difference.
Nonetheless, writing systems still play a role, especially for the $n$-gram-based models. Although topics-base and topics-translit achieve the same within-language accuracy (Tables 2/9), topics-translit performs slightly better cross-lingually. Its within-dataset performance is slightly lower than for topics-base (69.4% vs. 70.3%), likely due to some language-specific information being lost when diacritics are removed. This would be compensated in the within-language performance by improved transfer between datasets of languages that originally had different scripts.
E.3.1 Effects of transliteration
@@
Eight of the languages in SIB-200 come in two script versions each. The writing systems are completely distinct, except for Mandarin Chinese, which has versions in traditional and simplified characters. For topics-base, transfer between the two scripts of a language is unsurprisingly always much lower than the performance on the same script (with accuracies between 8.3% and 25.9% for in-language cross-script transfer – excluding the Mandarin Chinese entries, for which cross-script accuracy is up to 64.7% – vs. within-language, within-script accuracies between 60.8% and 75.9% for the same languages). Although transfer between the transliterated versions of the datasets works better (between 14.2% and 50.5%), the accuracies are still much lower than the within-language, within-original-script accuracies (between 58.3% and 74.0%). This is likely due to different transliteration conventions for different writing systems (and due to missing vowels in the transliterated versions of abjads).
The patterns are also similar for topics-mbert, with within-language, cross-script scores between 8.3% and 45.6%, compared to within-language, within-script accuracies between 42.2% and 79.9%.
@@ F.1 Correlations between task results
Table 10 shows the correlations between task results. It is described in §4.2. The correlations across different task types only involve the 55 training and 84 test languages that appear both in UD and in SIB-200 (54 and 82 in the comparisons with the transliterated SIB-200 data).
@@ F.3 Correlations between NLP results and similarity measures
Tables 14 and 15 show the correlations between POS/parsing scores and the similarity measures for each test language. Tables 16, 17, and 18 show the same for the topic classification experiments. Because these tables take up a lot of space, they are placed at the very end of the appendix.
Appendix G Implications for Choosing Source Languages
@@
This section provides additional context for §5.2.1. Table 12 shows the mean performance loss when choosing source languages based on their similarity to the target language (or based on their training data size). Table 12 shows the loss when choosing the source languages that were the best sources in another cross-lingual transfer task.